Project

General

Profile

Actions

20170905 » History » Revision 3

« Previous | Revision 3/5 (diff) | Next »
MacWaters, Craig , 06 September 2017 12:15


Ops/Comm Scheduling meeting : 5/9/2017

Location : CR21 at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Give LH2 documentation to Steve to add to official documentation list (Craig, Josef): completed draft exists but further changes expected; stands
    • Update shift checklist to focus on alarms : (Steve): stands (need to sit down with Paolo)
    • Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
  • DC update
  • Hydrogen system update and plans (JBoehm)
  • Status of alarm clearance (DC/Ajit)
  • Update on tracker readout (Melissa)
  • Update on DS Shield pressure (Henry)
  • Update on SS QPS connector issue (Craig)
  • Cycle 2017/01 Issues
  • Hall/Operations plan : attachment:17-08-17-CW.xlsx
  • Operations startup
    • Initial discussion on system readiness : Wednesday August 23rd in CR25 in R1 @ 11
    • Pre-startup meeting : Monday Sep 4th (Postponed until after the pulser run)
    • Pulser run : Thursday 7th : test of DAQ and online. Paolo will oversee this starting with a phone discussion
      to plan the day on Wednesday @ 16:00
  • Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
    No.
  • AOB

Minutes

Actions
Give LH2 SHIFTER documentation to Steve to add to official documentation list (Josef): completed draft exists but further changes expected from Josef & Andy; stands
Arrange a geometry survey for September just before running starts (Colin/ChrisR) : Tony Millington generally available, fix date closer to the time; stands
DC update
Hydrogen system update and plans (JBoehm)
Vent line leak checking in progress
Safety tour went well with only a few minor outstanding issues. Completed sign-off is expected in the next few days.
When leak checking complete, the aim is too cool with He over the weekend and then start with H2 first thing Monday.
Status of alarm clearance (DC/Ajit) - No update
Hall/Operations plan : 17-08-17-CW.xlsx
Operations startup
Pulser run : Sept 7th. Melissa, Agit and Durga to attend.
Beam tuning and beam test run : sometime in the week starting Sep 11th. Will ask ISIS for 8 hours during this week.
CRogers reported that ISIS will run at risk at 800MeV and could be susceptible to dipole problem. If they have to reduce to 700MeV then it will take a few days to sort out.
Update on Financial Restrictions In Technology and Engineering Subjects (CWhyte)
No.
AOB
DS is ok but unstable. Shield supply pressure is wavering. Mark T to investigate when freed from LH2 responsibilities

Updated by MacWaters, Craig about 6 years ago · 3 revisions